OpenStudy (anonymous):

log(sinxcosx)sinx*log(sinxcosx)cosx=1/4 sinxcosx is the base.

OpenStudy (turingtest):

\[\large \log_{\sin x\cos x}(\sin x)\log_{\sin x\cos x}(\cos x)=\frac14\sin x \cos x\]just wanted to type it more clearly so I can think about it as I wake up...

OpenStudy (anonymous):

After 1/4 there's no sinxcosx.

OpenStudy (turingtest):

\[\large\log_{\sin x\cos x}(\sin x)\log_{\sin x\cos x}(\cos x)=\frac14\]good thing I typed it then :)

OpenStudy (anonymous):

Sure is. :)

OpenStudy (anonymous):

\[\frac{1}{\log_{\sin x}\sin x + \log_{\sin x}\cos x} \times\frac{1}{\log_{\cos x}\sin x + \log_{\cos x}\cos x} \] \[\frac{1}{1 + \log_{\sin x}{\cos x}} \times \frac{1}{1 + \frac 1 {\log_{\sin x}\cos x}}\] \[\log_{\sin x}\cos x = t\] \[4t = (1 + t)^2\]

OpenStudy (anonymous):

\[(t - 1)^2 = 0\] \[t = 1\] \[\log_{\sin x}\cos x = 1\] \[\cos x = \sin x\] \[x = \frac{\pi}{4}\]You can get other solutions as well, but I am lazy :/

sam (.sam.):

The other solution is \[x=\frac{5\pi}{4}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're Welcome!